On heating 1.5 g of a mixture of kclo3 and kcl decreased to 0.12 g. If heating klolo3 creates kcl, calculate the percentage amount of kclo3 in the mixture.
1
Expert's answer
2020-08-17T13:33:14-0400

2KClO3 = 2KCl + 3O2

The mass of oxyged released 0.12 g

n(O2)= 0.12/32=0.00375 mol

n(KClO3)=2/3 n(O2) = 0.0025 mol

m(KClO3) = 0.30625 g

w(KClO3) = 0.30625 / 1.5 = 20.42%

Answer: w(KClO3) = 20.42%